OpenStudy (anonymous):

(1/12)+(3/8)y=(5/12)+(5/8)y

OpenStudy (anonymous):

Well, I would start out by combining like terms: Subtract 1/12 from both sides and 5y/8 from both sides and see what you come up with.

OpenStudy (anonymous):

What do you wan to find exactly? If you mean is y, here my answer 1/12 - 5/12 = 5/8 y - 3/8 y -4 / 12 = 1/4 y y = -3/4